OpenStudy (hippoharry):

The equation 2x^2 - 2x - 12 = 0 is factored. Each factor is set equal to zero. What are these two equations? A. 2x - 6 = 0, x + 2 = 0 B. 2x - 3 = 0, x + 4 = 0 C. 2x - 4 = 0, x + 3 = 0 D. 2x + 6 = 0, x - 2 = 0

OpenStudy (superpanda1022):

A

OpenStudy (superpanda1022):

Because (2x-6)(x+2)=2x^2-2x-12

OpenStudy (calculusxy):

Using the factor by grouping method you would do the following steps: 2x^2 - 2x - 12 = 0 2x^2 + 4x - 6x - 12 2x(x + 2) -6(x+2) (2x-6)(x+2) = 0 So to find the roots, you would need to set each of them equal to zero. 2x - 6 = 0 ; x + 2 = 0